# Segment Anything Model for Road Network Graph Extraction

### v2

@@ 1 Introduction
Recently, the rapid growth of foundational models [2, 45, 52, 46] showcased their impressive capabilities. These models, which leverage flexible, high-capacity, and scalable architectures such as Transformers [53], are pre-trained through effective self-supervision [22] methods and unprecedentedly large datasets. This endows them with robust semantic reasoning and generalization. Segment Anything Model (SAM) [29] is such a foundational vision model. Built on the powerful ViT [16] architecture and trained with billions of images and masks, it demonstrates unparalleled semantic segmentation capabilities. This raises intriguing questions: How can SAM be applied to the prediction of road network graphs from satellite images, and how good can it be?
@@ 2.1 SAM and Its Applications
In 2023, the Segment Anything paper [29] proposed a foundational model for image segmentation, capable of generalizing across various tasks, enabling impressive zero-shot transfer capabilities. Through fine-tuning or direct adoption, SAM has been used in object detection [36], image inpainting [62], segmentation of medical images [39, 28, 55, 64], and remote sensing tasks [11]. Existing adaptations of SAM in remote sensing have focused more on simple segmentation and have not yet been applied to the production of road network graphs.
@@ 3.3 Geometry Decoder
After acquiring the masks, the graph vertices are extracted from them. This process converts the dense mask images into a set of sparse vertices, with roughly the same interval $d_{v}$ in between. It’s implemented with simple non-maximum suppression: we first drop the pixels under a probability threshold $t$, then traverse them by a descending order of their probability. Pixels within a $d_{v}$ radius of the current one are removed. The $(x,y)$ locations of the remaining pixels form the graph vertices $\mathbf{V}\{v_{i}\in\mathbb{R}^{2}\}$ See Algorithm 1.
@@
1:$\mathbf{V}\leftarrow\emptyset$triangleright Initialize the set of pixel coordinates
@@
We predict masks for both intersections and roads for more accurate graph structures at intersections. If only the road mask existed, there would be no guarantee that the center point of an intersection would be kept, producing error patterns like Figure  6. To mitigate this: 1) Vertices are extracted from both masks with the same NMS algorithm. 2) The two sets of vertices are joined, with all intersection vertices assigned a higher score than any road vertices. 3) The joined set is then NMS-ed again to produce the final result. This ensures intersection points are kept as much as possible.
@@ 3.4 Topology Decoder
The connection here is defined as ”whether two vertices are immediate neighbors on the graph”. That is, imagine a breadth-first-search on the road network graph from the source vertex, which stops expanding whenever a) it hits a target vertex or b) the depth (search radius) exceeds $R_{\text{nbr}}$ a target vertex is only connected to the source if it’s visited by the search. This is further illustrated in Figure  3.
@@ 4.3 Implementation Details
For both datasets, $R_{\text{nbr}}$ is 64 pixels (meters), $N_{\text{nbr}}$ is 16. At training time, For City-scale, we sample image patches of $512\times 512$ pixel, the batch size is 16 and we sample 512 source points for topology query per image patch. For SpaceNet, the batch size is 64 due to using image patches $256\times 256$ pixel. We sample 128 source points per patch. When there are fewer than $N_{\text{nbr}}$ available target nodes to query, we use attention masking to ensure the interaction only happens between the valid vertices.
@@ 4.4 Evaluating Road Network Prediction
Table 1: Comparison with existing methods on different datasets. SAM-Road achieved the highest TOPO precision of $90.47\%$ on City-scale and $93.03\%$ on SpaceNet. It also shows the highest APLS metric of on both sets. Overall the graph accuracy is among the very top.
Qualitative results of SAM-Road predicting large-scale road network graphs can be found in Figure  5. The results are shown side-by-side with two baselines and the ground-truths. Overall, SAM-Road predicts highly accurate road networks even under very challenging circumstances, e.g. many blocks and intersections in dense urban areas, curvy roads with irregular shapes, overpasses, and multi-lane highways.
@@
Such performance should largely be attributed to SAM features, the highly powerful foundational vision model. As shown in Figure 1, the predicted masks are sharp and clear, enabling precise geometry prediction. The SAM image features are also informative vertex embeddings containing rich semantic meanings, as evident in the accurate topology predictions.
@@ 4.5 Speed and Accuracy Trade-off
SAM-Road is also highly efficient, thanks to its parallelized inference and it doesn’t require complex CPU-heavy post-processing heuristics. We measure the inference time to produce the complete graphs for the test sets of both datasets. The main results use $16\times 16$ windows and are already 40×40$\times$40 × faster than RNGDet++ on the City-scale dataset, and 10×10$\times$10 × faster on the SpaceNet dataset, as shown in Table 2. As mentioned in Section 3.6, SAM-Road can trade accuracy for more speed by sparsifying the sliding windows. Table 3 shows the result such trade-off. Using fewer windows can further provide 2×2$\times$2 × to 4×4$\times$4 × speed-up, with a minor accuracy drop.
@@ 4.6 Ablation Studies
Whether using the learning-based topology decoder. A vs G shows that it’s critical for SAM-Road’s performance. Intuitively, a naive method that might achieve a similar effect is just to run a pathfinding algorithm between a pair of vertices, using the road existence map as the cost field, and see if there’s a sufficiently low-cost path between the two without passing through other vertices. We implemented such a variant G using an A-star algorithm. Metric is much worse, and qualitatively as shown in Figure  6. This approach can mess up intersections, overpasses, and close parallel roads.
@@ 5 Limitations and Future Work
One current limitation of SAM-Road is we have not designed specific approaches to more accurately handle overpasses. There is an ambiguity for the topology decoder at the exact point where overpassing roads intersect. However, this issue is minor as most vertices are not at these crossing points. Future work could improve this by predicting an overpass heatmap to suppress vertex formation at these locations.
In addition, in this work, we only utilized the smallest Segment Anything model, ViT-B. Larger variants may be explored as a future work, where we hope to explore parameter-efficient tuning methods, such as LoRA [27].
We are also interested in exploring the integration of state-of-the-art Vision Language Models (VLMs) such as PaLI [12] and GPT-4V [1] with graph learning.

### v3

@@ 1 Introduction
Recently, the rapid growth of foundational models [2, 46, 53, 47] showcased their impressive capabilities. These models, which leverage flexible, high-capacity, and scalable architectures such as Transformers [54], are pre-trained through effective self-supervision [22] methods and unprecedentedly large datasets. This endows them with robust semantic reasoning and generalization. Segment Anything Model (SAM) [29] is such a foundational vision model. Trained with millions of images and billions of masks, it demonstrates unparalleled semantic segmentation capabilities. This raises intriguing questions: How can SAM be applied to the prediction of road network graphs from satellite images, and how good can it be?
@@ 2.1 SAM and Its Applications
In 2023, Segment Anything Model [29] was proposed as a foundational model for image segmentation, showcasing impressive zero-shot and generalization capabilities. Through fine-tuning or direct adoption, SAM has been used in object detection [36], image inpainting [63], segmentation of medical images [39, 28, 56, 65], and remote sensing tasks [11]. Existing adaptations of SAM in remote sensing have focused more on simple segmentation and have not yet been applied to the production of road network graphs.
@@ 3.3 Geometry Decoder
After acquiring the masks, the graph vertices are extracted from them. This process converts the dense mask images into a set of sparse vertices, with roughly the same interval $d_{v}$ in between. $d_{v}$ is selected to be sparse while not hurting geometry accuracy. It’s implemented with simple non-maximum suppression: we first drop the pixels under a probability threshold $t$, then traverse them by a descending order of their probability. Pixels within a $d_{v}$ radius of the current one are removed. The $(x,y)$ locations of the remaining pixels form the graph vertices $\mathbf{V}\{v_{i}\in\mathbb{R}^{2}\}$ See Algorithm 1.
@@
1:$\mathbf{V}\leftarrow\emptyset$
@@
We predict masks for both intersections and roads for more accurate graph structures at intersections. If only the road mask existed, there would be no guarantee that the center point of an intersection would be kept, producing error patterns like Figure  6. To mitigate this: 1) Vertices are extracted from both masks with the same NMS algorithm. 2) The two sets of vertices are joined, with all intersection vertices assigned a higher score than any road vertices. 3) The joined set is then NMS-processed again to produce the final result. This ensures intersection points are kept as much as possible.
@@ 3.4 Topology Decoder
The connection here is defined as ”whether two vertices are immediate neighbors on the graph”. That is, imagine a breadth-first-search on the road network graph from the source vertex, which stops expanding whenever a) it hits a target vertex or b) the depth (search radius) exceeds $R_{\text{nbr}}$ a target vertex is only connected to the source if it is visited by the search. This is further illustrated in Figure  3.
@@ 4.3 Implementation Details
For both datasets, $d_{v}$ is 16 pixels (meters), $R_{\text{nbr}}$ is 64 pixels (meters), $N_{\text{nbr}}$ is 16, $D_{\text{feat}}$ is 128. At training time, For City-scale, we sample image patches of $512\times 512$ pixel, the batch size is 16 and we sample 512 source points for topology query per image patch. For SpaceNet, the batch size is 64 due to using image patches $256\times 256$ pixel. We sample 128 source points per patch. When there are fewer than $N_{\text{nbr}}$ available target nodes to query, we use attention masking to ensure the interaction only happens between the valid vertices.
@@ 4.4 Evaluating Road Network Prediction
Table 1: Comparison with existing methods on different datasets. SAM-Road achieved the highest TOPO precision of $90.47\%$ on City-scale and $93.03\%$ on SpaceNet. It also shows the highest APLS metric of on both sets. Overall the graph accuracy is among the very top. SAM-Road leans more towards precision in TOPO metrics, this might be due to the low positive / negative example ratio in its binary classification tasks.
Qualitative results of SAM-Road predicting large-scale road network graphs can be found in Figure  5. The results are shown side-by-side with two baselines and the ground-truths. Some error examples can be found in Figure 7. Overall, SAM-Road predicts highly accurate road networks even under very challenging circumstances, e.g. many blocks and intersections in dense urban areas, curvy roads with irregular shapes, overpasses, and multi-lane highways.
@@
Such performance should largely be attributed to SAM, the powerful foundational vision model. As shown in Figure 1, the predicted masks are sharp and clear, enabling precise geometry prediction. The SAM image features are also informative vertex embeddings containing rich semantic meanings, as evident in the accurate topology predictions.
@@ 4.5 Speed and Accuracy Trade-off
SAM-Road is also highly efficient, thanks to its parallelized inference and that it doesn’t require complex CPU-heavy post-processing heuristics. We measure the inference time to produce the complete graphs for the test sets of both datasets. The main results use $16\times 16$ windows and are already 40×40$\times$40 × faster than RNGDet++ on the City-scale dataset, and 10×10$\times$10 × faster on the SpaceNet dataset, as shown in Table 2. As mentioned in Section 3.6, SAM-Road can trade accuracy for more speed by sparsifying the sliding windows. Table 3 shows the result such trade-off. Using fewer windows can further provide 2×2$\times$2 × to 4×4$\times$4 × speed-up, with a minor accuracy drop.
@@ 4.6 Ablation Studies
Whether using the learning-based topology decoder. A vs G shows that it’s critical for SAM-Road’s performance. Intuitively, a naive method that might achieve a similar effect is just to run a pathfinding algorithm between a pair of vertices, using the road existence map as the cost field, and see if there’s a sufficiently low-cost path between the two without passing through other vertices. We implemented such a variant G using an A-star algorithm. Metrics are much worse, as qualitatively shown in Figure  6. This approach can mess up intersections, overpasses, and close parallel roads.
@@ 5 Limitations and Future Work
Figure 7: Some error patterns. Left: geometry decoder missed the road segment in the middle. Middle: topology decoder missed connections in a complex interchange. Right: an interesting case where SAM-road predicts the trails in a park which are not part of the label.
One current limitation of SAM-Road is we have not designed specific approaches to more accurately handle overpasses. There is an ambiguity for the topology decoder at the exact point where overpassing roads intersect, as the correct answer depends on which layer is being asked. This issue is minor though, as most vertices are not at these spots. Future work could improve this by predicting an overpass heatmap to suppress vertex formation at these locations.
In addition, in this work, we only used the smallest Segment Anything model, ViT-B. Larger variants may be explored as a future work, where we hope to explore parameter-efficient tuning methods, such as LoRA [27].
We are also interested in exploring the integration of other state-of-the-art foundational models, such as DINOv2 [44], PaLI [12] and GPT-4V [1] with graph learning.